If the volume of H2 produced in the experiment is 73.0 ml, the pressure is 745.6 torr, the temperature is 25.5 °C, moles of H2 is 0.003, what is the gas constant R in the unit of L•atm/mol•K?
Answer #1

Given:
V = 73.0 mL
= 0.0730 L

P = 745.6 torr
= 745.6 / 760 atm
= 0.9811 atm
T = 25.5 oc
= (25.5 + 273.1) K
= 298.6 K
n = 0.003 mol

Use:
P*V = n*R*T
0.9811 atm * 0.0730 L = 0.003 mol * R * 298.6 K
R = 0.07995 L.atm/mol.K

Answer: 0.080